# Multimodal Point-of-Interest Recommendation

### v1

@@ 2.1.1. Base Datasets
We choose the Foursquare dataset and FoodX-251 dataset (Table 1). The Foursquare dataset contains check-ins in New York City, United States and Tokyo, Japan from 12 April 2012 to 16 February 2013, and we used check-ins in Tokyo (573,703 in total). Each check-in has a user ID, a time stamp, a set of geographical coordinates (latitude and longitude), a venue ID, a venue category (a pair of a name and an ID). The venue categories are very detailed and we select food-related ones manually (2.1.2).
The FoodX-251 dataset contains 158,846 food images collected from the web and labelled in 251 classes. We select 106 classes based on check-in frequency in corresponding Foursquare’s venue category (2.1.4).
@@ 2.1.2. Meta Data and Interaction Data
To extract check-ins at food-related POIs, we select 80 out of 247 categories defined in the Foursquare dataset (Appendix A), and there are 31,105 POIs under these categories. We finally get 28,989 POIs with attributes as explained later (we call this attribute dataset as meta data.) We extract check-ins at these POIs from the Foursquare dataset, filter check-ins by loyal users with 100 or more check-ins and get 119,105 check-ins by 2,092 users (we call this check-in dataset as interaction data.)
@@ 3.2. Recformer
As explained in 3.1, our input to the Recformer model is a sequence of truncated attribute texts of each venue. We conduct pretraining with Masked Language Modelling (Devlin et al., 2019) and item-item contrastive task which are widely used to predict the next item to recommend, and then take Two-Stage Finetuning (Li et al., 2023) with item-item contrastive learning as learning task. We keep the task settings of the previous work (Li et al., 2023).
@@ 4.1.3. Inference
As a preparation, we create a list of embedded values of all the POIs using the model trained in 4.1.2.
After that, we take 3 steps for each sequence in the test dataset to predict the next POI: First, we keep the last POI of the sequence for performance evaluation (4.1.4) and use the rest of the sequence. Then, we put this part of the sequence (except the last POI) into the trained model and get an embedded value of the next POI (prediction of the last POI). Finally, we select the most similar POI to this value from the list of embedded POIs in terms of cosine similarity.

### v2

@@ 2.1.1. Base Datasets
We choose the Foursquare dataset and FoodX-251 dataset (Table 1). The Foursquare dataset contains check-ins in New York City, United States and Tokyo, Japan from 12 April 2012 to 16 February 2013, and we used check-ins in Tokyo (573,703 in total). Each check-in has a user ID, a time stamp, a set of geographical coordinates (latitude and longitude), a venue ID, a venue category (a pair of a name and an ID). The venue categories are very detailed and we select food-related ones manually (2.1.2). The FoodX-251 dataset contains 158,846 food images collected from the web and labelled in 251 classes. We select 106 classes based on check-in frequency in corresponding Foursquare’s venue category (2.1.4).
@@ 2.1.2. Meta Data and Interaction Data
To extract check-ins at food-related POIs, we select 80 out of 247 categories defined in the Foursquare dataset (Appendix A), and there are 31,105 POIs under these categories. We finally get 28,989 POIs with attributes as explained later (we call this attribute dataset as meta data.) We extract check-ins at these POIs from the Foursquare dataset, filter check-ins by loyal users with 100 or more check-ins and get 119,105 check-ins by 2,092 users.
@@ 3.2. Recformer
As explained in 3.1, the model input is a sequence of truncated attribute texts of each venue. We conduct pretraining with Masked Language Modelling (Devlin et al., 2019) and item-item contrastive task which are widely used to predict the next item to recommend, and then take Two-Stage Finetuning (Li et al., 2023) with item-item contrastive learning as learning task. We keep the task settings of the previous work (Li et al., 2023).
@@ 4.1.3. Inference
As a preparation, we create a list of embedded values of all the POIs using the model trained in 4.1.2. Then, we take 3 steps for each sequence in the test dataset to predict the next POI: First, we keep the last POI of the sequence for performance evaluation (4.1.4) and use the rest of the sequence. Then, we put this part of the sequence (except the last POI) into the trained model and get an embedded value of the next POI (prediction of the last POI). Finally, we select the most similar POI to this value from the list of embedded POIs in terms of cosine similarity.
@@ 5. Conclusion and Future Work
Acknowledgements.
This work is partially supported by ”Joint Usage/Research Center for Interdisciplinary Large-scale Information Infrastructures (JHPCN)” in Japan (Project IDs: jh231004 and jh241004). In addition, this work was partially supported by JSPS KAKENHI Grant Numbers JP21K17749 and 23K28098.